side a (0,0) to (3,0)
side b (3,0) to (3,4)
find the length of c:

Answer by rfer(16322)

3^2+4^2=c^2
9+16=c^2
25=c^2
c=sqrt25
c=5
-------
By the way this is a standard 3, 4, 5 triangle